OpenStudy (anonymous):

How many solutions (using only nonnegative integers) are there to the following equation? x1 + x2 + x3 + x4 + x5 = 36

OpenStudy (unklerhaukus):

so, n = 5 r = 36

OpenStudy (unklerhaukus):

C(n + r-1, r) =

OpenStudy (anonymous):

I got this far this is what i have. c(40, 36) c(40*39*38*37)(4*3*2*1) from here on out I don't know how to solve it

OpenStudy (unklerhaukus):

C(40, 36) = (40*39*38*37) / (4*3*2*1) =

OpenStudy (unklerhaukus):

now, cancel common factors,

OpenStudy (unklerhaukus):

= 10 * 13 * 19 * 37

OpenStudy (unklerhaukus):

[ you can use a calculator now ]

OpenStudy (anonymous):

91390 Thanx it was the confusing part thanx